# V2M4: 4D Mesh Animation Reconstruction from a Single Monocular Video

### v1

@@ Abstract
We present V2M4, a novel 4D reconstruction method that directly generates a usable 4D mesh animation asset from a single monocular video. Unlike existing approaches that rely on priors from multi-view image and video generation models, our method is based on native 3D mesh generation models. Naively applying 3D mesh generation models to generate a mesh for each frame in a 4D task can lead to issues such as incorrect mesh poses, misalignment of mesh appearance, and inconsistencies in mesh geometry and texture maps. To address these problems, we propose a structured workflow that includes camera search and mesh reposing, condition embedding optimization for mesh appearance refinement, pairwise mesh registration for topology consistency, and global texture map optimization for texture consistency. Our method outputs high-quality 4D animated assets that are compatible with mainstream graphics and game software. Experimental results across a variety of animation types and motion amplitudes demonstrate the generalization and effectiveness of our method.
@@ 3 Preliminaries
Our method builds upon TRELLIS [49], a recently introduced and powerful 3D generation model. TRELLIS encodes 3D assets into a structured latent representation (SLAT), which consists of a sparse 3D grid embedding both geometric and appearance information. The 3D reconstruction process in TRELLIS can be broadly divided into two phases. In the first phase, TRELLIS employs rectified flow  [3] to generate a 3D voxel grid, where the voxel values indicate the object’s occupancies. By applying a threshold, TRELLIS extracts a sparse voxel representation. In the second phase, a sparse rectified flow transformer is used to generate the object’s SLAT from this sparse voxel grid. Both phases are conditioned on visual features extracted from the input image using DINOv2  [28], along with a negative condition based on the classifier-free technique  [9]. The SLAT can then be processed by three independent decoders, which generate meshes, Gaussian Splats  [18], and Radiance Fields  [6], respectively. For mesh decoding, TRELLIS predicts both the signed distance field (SDF) grid and the parameters of differentiable FlexiCubes  [36], enabling the conversion from the SDF into a mesh. Instead of using the predicted mesh vertex colors from the mesh decoder, TRELLIS employs xatlas  [56] to compute UV maps of the mesh and then bakes a high-resolution texture map using multi-view observations of its Gaussian Splats.
@@ 4.1 Camera Search and Mesh Re-Pose
Figure 3: Camera Pose Extraction from Dense Stereo Reconstruction for $V_{\text{ref},t}$. We first obtain estimated point clouds $\overline{\mathrm{PC}}_{\text{ref}}$ and $\overline{\mathrm{PC}}_{\text{rend}}$ from the DUSt3R model for reference video frame $V_{\text{ref},t}$ and rendering views $V_{\text{rend}}$. We then align $\overline{\mathrm{PC}}_{\text{rend}}$ with the point clouds ${\mathrm{PC}}_{\text{rend}}$ obtained from mesh ~$\widetilde{\mathcal{M}}_{t}$ during rasterization (Note the differences in axis direction and scale depicted in the figure.). This transformation is then applied to $\overline{\mathrm{PC}}_{\text{ref}}$, resulting in the transformed ${\mathrm{PC}}_{\text{ref}}$. Finally, we optimize to determine the camera pose $\mathcal{C}_{\text{ref},t}$ that ensures accurate mapping from 3D points to 2D points.
@@
Given an approximately aligned camera pose where the rendering view matches $V_{\text{ref},t}$., we further refine this alignment by applying gradient descent optimization on the camera pose. Specifically, we minimize the discrepancy between the mask region of the rendering view—defined as the pixels covered by the object during rasterization—and the object region in $V_{\text{ref},t}$. With a well-initialized starting point, the gradient descent efficiently converges to a near-optimal solution for $\mathcal{C}_{\text{ref},t}$.
@@ 4.3 Consistent Topology via Iterative Pairwise Registration
Due to the inherent randomness of the generation model, reconstructed meshes from adjacent video frames often exhibit variations in topology, such as differing numbers of vertices and faces, as well as changes in vertex and edge connectivity. To create a valid 4D animation asset with consistent geometry, we propose an iterative pairwise registration approach. For two consecutive meshes after refinement, ~$\widetilde{\mathcal{M}}_{\text{refine},t}$ and ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$, we begin by treating ~$\widetilde{\mathcal{M}}_{\text{refine},t}$ as a rigid body. We optimize its global transformation (rotation, translation, and scale) to align it with ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$ using Chamfer distance and a differentiable rendering loss. After achieving global registration, we refine the alignment using the preconditioned optimization proposed by  [27], which allows for fast and smooth convergence in the inverse reconstruction process. However, the differentiable rendering loss from  [27] can struggle with regions undergoing large deformations or containing long, thin structures. To improve registration in these areas, we incorporate the As-Rigid-As-Possible (ARAP) constraint  [13] alongside the Chamfer distance for enhanced local registration. Through these global and local alignment steps, we obtain a deformed version ~$\widetilde{\mathcal{M}}^{\prime}_{t}$ (with vertices $\mathcal{V}^{\prime}_{t}$ and faces $\mathcal{F}^{\prime}_{t}$) of ~$\widetilde{\mathcal{M}}_{\text{refine},t}$ ($\mathcal{V}_{t}$ and faces $\mathcal{F}_{t}$), ensuring that the number and topology are identical ($\mathcal{V}^{\prime}_{t}=\mathcal{V}_{t}$, $\mathcal{F}^{\prime}_{t}=\mathcal{F}_{t}$) and closely resemble the shape of ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$. We then replace the initial ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$ with the deformed ~$\widetilde{\mathcal{M}}^{\prime}_{t}$. In practice, we designate ~$\widetilde{\mathcal{M}}_{\text{refine},1}$ as the rest pose and iteratively apply the registration process across consecutive frames. This ensures that all meshes ~$\widetilde{\mathcal{M}}^{\prime}_{\{1,\dots,T\}}$ maintain consistent geometry.
@@ 4.5 Mesh Interpolation and 4D Asset Conversion
Table 2: Quantitative Evaluation of Different Methods. All metrics are computed using the rendering view of the reconstructed meshes. “Simple” refers to the Consistent4D data, which consists of simple objects with subtle movements, while “Complex” refers to data collected online featuring more complex objects and large-scale movements. “All” represents the combined evaluation on both subsets. The best results are highlighted in bold.
@@
Table 3: Average Reconstruction Time per Frame. The displayed time represents the average time taken to reconstruct each frame across an animation sequence.
@@ 5.1 Experimental Setup
Please refer to Appendix A for implementation details, and Appendix B&C for ablation studies and more visual results.
@@
We constructed a benchmark comprising 40 animation videos. Specifically, we first collected 20 animation videos from Consistent4D [15], which feature simple object topologies with subtle movements. Additionally, we gathered 20 animation videos from Mixamo [25] and Sketchfab [39] to include larger-scale object movements and more complex object topologies, thus providing a diverse test set for assessing reconstruction performance. Our benchmark is much larger and more varied than those used in existing works [20, 15, 57], which evaluate only 7 test samples. Sample frames from our benchmark can be found in Appendix C. Since existing video segmentation methods [30] can easily extract foreground object regions, the dataset has been pre-processed to remove backgrounds. This ensures that the evaluation remains focused on reconstruction quality rather than segmentation challenges.
@@ 5.3 Limitations
While our method effectively reconstructs 4D mesh animations from input videos in many scenarios, it has several limitations. First, since our approach is based on TRELLIS, our method may suffer from performance issues when TRELLIS outputs poor 3D reconstructions. Second, artifacts can occur when reconstructing animations with topology changes, as we enforce topology consistency by registering the first mesh to subsequent frames. Third, using a fixed sampling rate, as discussed in Section 4.5, can be inefficient due to variable motion dynamics—sometimes moving slowly and other times rapidly. Future work could benefit from incorporating adaptive sampling techniques.
@@ Appendix A Implementation Details
In the utilization of the dense stereo reconstruction model, we obtain the ground truth point clouds of ~$\widetilde{\mathcal{M}}_{t}$ during the rasterization process using Nvdiffrast [19]. To speed up optimization and filter out possible outliers in the point clouds, we retain only a 5% points of both the ground truth point clouds $\mathrm{PC}_{\text{rend},{\{1,\dots,n\}}}$ and the predicted ones $\overline{\mathrm{PC}}_{\text{rend},{\{1,\dots,n\}}}$. We then align these two sets of point clouds using Chamfer loss to optimize the global transformation (rotation, scale, and translation) over 2000 iterations. To prevent the alignment from becoming trapped in local minima, such as an object being flipped along its vertical axis, we manually flip the point clouds along the axis starting at the 1000th iteration. We then compare the Chamfer loss before and after flipping to select the best alignment. After aligning the point cloud sets, we optimize the camera pose for 500 iterations using MSE loss between the 3D point clouds of $\mathrm{PC}_{\text{ref}}$ and the 2D pixel positions in $V_{\text{ref},t}$.
@@
By computing the visual alignment between the rendering view of the reconstructed mesh and Gaussian Splats, we backpropagate the gradient through differentiable rendering to the negative condition embedding in TRELLIS. We use a combination of DreamSim, LPIPS, and MSE losses (each with a weight of 1), along with an MSE regularization term weighted at 0.2 to compare with the embedding’s initial value, preventing distortion of the reconstructed result. This optimization is applied only during the latter part of the flow model ($0.6\geq t\geq 0$, where $t$ represents the timestep of the flow model). We begin the optimization at iteration 5 and gradually increase by 1 every 5 timesteps of the flow model.

### v2

@@ Abstract
We present V2M4, a novel 4D reconstruction method that directly generates a usable 4D mesh animation asset from a single monocular video. Unlike existing approaches that rely on priors from multi-view image and video generation models, our method is based on native 3D mesh generation models. Naively applying 3D mesh generation models to generate a mesh for each frame in a 4D task can lead to issues such as incorrect mesh poses, misalignment of mesh appearance, and inconsistencies in mesh geometry and texture maps. To address these problems, we propose a structured workflow that includes camera search and mesh reposing, condition embedding optimization for mesh appearance refinement, pairwise mesh registration for topology consistency, and global texture map optimization for texture consistency. Our method outputs high-quality 4D animated assets that are compatible with mainstream graphics and game software. Experimental results across a variety of animation types and motion amplitudes demonstrate the generalization and effectiveness of our method. https://windvchen.github.io/V2M4.
@@ 3 Preliminaries
Our method builds upon TRELLIS [52], a recently introduced and powerful 3D generation model (we also evaluate our method using more advanced 3D generators, as detailed in Appendix B). TRELLIS encodes 3D assets into a structured latent representation (SLAT), which consists of a sparse 3D grid embedding both geometric and appearance information. The 3D reconstruction process in TRELLIS can be broadly divided into two phases. In the first phase, TRELLIS employs rectified flow  [3] to generate a 3D voxel grid, where the voxel values indicate the object’s occupancies. By applying a threshold, TRELLIS extracts a sparse voxel representation. In the second phase, a sparse rectified flow transformer is used to generate the object’s SLAT from this sparse voxel grid. Both phases are conditioned on visual features extracted from the input image using DINOv2  [30], along with a negative condition based on the classifier-free technique  [9]. The SLAT can then be processed by three independent decoders, which generate meshes, Gaussian Splats  [18], and Radiance Fields  [6], respectively. For mesh decoding, TRELLIS predicts both the signed distance field (SDF) grid and the parameters of differentiable FlexiCubes  [38], enabling the conversion from the SDF into a mesh. Instead of using the predicted mesh vertex colors from the mesh decoder, TRELLIS employs xatlas  [59] to compute UV maps of the mesh and then bakes a high-resolution texture map using multi-view observations of its Gaussian Splats.
@@ 4.1 Camera Search and Mesh Re-Pose
Figure 3: Camera Pose Extraction from Dense Stereo Reconstruction for $V_{\text{ref},t}$. We first obtain estimated point clouds $\overline{\mathrm{PC}}_{\text{ref}}$ and $\overline{\mathrm{PC}}_{\text{rend}}$ from the DUSt3R model for reference video frame $V_{\text{ref},t}$ and rendering views $V_{\text{rend}}$. We then align $\overline{\mathrm{PC}}_{\text{rend}}$ with the point clouds ${\mathrm{PC}}_{\text{rend}}$ obtained from mesh ~$\widetilde{\mathcal{M}}_{t}$ during rasterization (Note the differences in axis direction and scale depicted in the figure.). This transformation is then applied to $\overline{\mathrm{PC}}_{\text{ref}}$, resulting in the transformed ${\mathrm{PC}}_{\text{ref}}$. Finally, we optimize to determine the camera pose $\mathcal{C}_{\text{ref},t}$ that ensures accurate mapping from 3D to 2D points.
@@
Given an approximately aligned camera pose where the rendering view matches $V_{\text{ref},t}$, we further refine this alignment by applying gradient descent optimization on the camera pose. Specifically, we minimize the discrepancy between the mask region of the rendering view—defined as the pixels covered by the object during rasterization—and the object region in $V_{\text{ref},t}$. With a well-initialized starting point, the gradient descent efficiently converges to a near-optimal solution for $\mathcal{C}_{\text{ref},t}$.
@@ 4.3 Consistent Topology via Iterative Pairwise Registration
Due to the inherent randomness of the generation model, reconstructed meshes from adjacent video frames often exhibit variations in topology, such as differing numbers of vertices and faces, as well as changes in vertex and edge connectivity. To create a valid 4D animation asset with consistent geometry, we propose an iterative pairwise registration approach. For two consecutive meshes after refinement, ~$\widetilde{\mathcal{M}}_{\text{refine},t}$ and ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$, we begin by treating ~$\widetilde{\mathcal{M}}_{\text{refine},t}$ as a rigid body. We optimize its global transformation (rotation, translation, and scale) to align it with ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$ using Chamfer distance and a differentiable rendering loss. After achieving global registration, we refine the alignment using the preconditioned optimization proposed by  [29], which allows for fast and smooth convergence in the inverse reconstruction process. However, the differentiable rendering loss from  [29] can struggle with regions undergoing large deformations or containing long, thin structures. To improve registration in these areas, we incorporate the As-Rigid-As-Possible (ARAP) constraint  [13] alongside the Chamfer distance for enhanced local registration. Through these global and local alignment steps, we obtain a deformed version ~$\widetilde{\mathcal{M}}^{\prime}_{t}$ (with vertices $\mathcal{V}^{\prime}_{t}$ and faces $\mathcal{F}^{\prime}_{t}$) of ~$\widetilde{\mathcal{M}}_{\text{refine},t}$ ($\mathcal{V}_{t}$ and faces $\mathcal{F}_{t}$), ensuring that the number and topology are identical ($\mathcal{V}^{\prime}_{t}=\mathcal{V}_{t}$, $\mathcal{F}^{\prime}_{t}=\mathcal{F}_{t}$) and closely resemble the shape of ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$. We then replace the initial ~$\widetilde{\mathcal{M}}_{\text{refine},t+1}$ with the deformed ~$\widetilde{\mathcal{M}}^{\prime}_{t}$. In practice, we designate ~$\widetilde{\mathcal{M}}_{\text{refine},1}$ as the rest pose and iteratively apply the registration process across consecutive frames. This ensures consistent geometry across all meshes ~$\widetilde{\mathcal{M}}^{\prime}_{\{1,\dots,T\}}$.
@@ 4.5 Mesh Interpolation and 4D Asset Conversion
Table 2: Quantitative Evaluation of Different Methods. All metrics are computed using the rendering view of the reconstructed meshes. “Simple” refers to the Consistent4D data, which consists of simple objects with subtle movements, while “Complex” refers to data collected online featuring more complex objects and large-scale movements. “All” represents the combined evaluation on both subsets. The best results are in bold.
@@
Table 3: Average Reconstruction Time per Frame. The displayed time represents the average time taken to reconstruct each frame across an animation sequence. Detailed timing for each step of our method is provided in Appendix A.
@@ 5.1 Experimental Setup
Please refer to Appendix A for implementation details, and to Appendix B & D for comprehensive ablation studies (including experiments with more advanced 3D generation models such as Hunyuan3D-2.0 [67], TripoSG [21], and advanced dense stereo reconstruction models like VGGT [48]), as well as additional visual results.
@@
We constructed a benchmark comprising 40 animation videos. Specifically, we first collected 20 animation videos from Consistent4D [15], which feature simple object topologies with subtle movements. Additionally, we gathered 20 animation videos from Mixamo [27] and Sketchfab [41] to include larger-scale object movements and more complex object topologies (with sequences containing more than 200 frames), thus providing a diverse test set for assessing reconstruction performance. Our benchmark is much larger and more varied than those used in existing works [22, 15, 60], which evaluate only 7 test samples (each with just 32 frames). Sample frames from our benchmark can be found in Appendix D. Since existing video segmentation methods [32] can easily extract foreground object regions, the dataset has been pre-processed to remove backgrounds. This ensures that the evaluation remains focused on reconstruction quality rather than segmentation challenges.
@@ 5.3 Limitations
While our method effectively reconstructs 4D mesh animations from input videos in many scenarios, it has several limitations. First, since our approach is based on TRELLIS, our method may suffer from performance issues when TRELLIS outputs poor 3D reconstructions (see further analysis in Appendix B). Second, artifacts can occur when reconstructing animations with topology changes, as we enforce topology consistency by registering the first mesh to subsequent frames. Third, using a fixed sampling rate, as discussed in Section 4.5, can be inefficient due to variable motion dynamics—sometimes moving slowly and other times rapidly. Future work could benefit from incorporating adaptive sampling techniques.
@@ 6 Conclusion
Acknowledgements
This work was supported by funding from King Abdullah University of Science and Technology (KAUST) — Center of Excellence for Generative AI, under award number 5940.
@@ Appendix A Implementation Details
In Sec. 4 of the main paper, we described the design of each component in our framework. These components are sequentially integrated according to the overall pipeline shown in Fig. 2 of the main paper. Below, we provide additional implementation details for each component of our framework.
@@
In the utilization of the dense stereo reconstruction model, we obtain the ground truth point clouds of ~$\widetilde{\mathcal{M}}_{t}$ during the rasterization process using Nvdiffrast [19]. To speed up optimization and filter out possible outliers in the point clouds, we retain only 5% points of both the ground truth point clouds $\mathrm{PC}_{\text{rend},{\{1,\dots,n\}}}$ and the predicted ones $\overline{\mathrm{PC}}_{\text{rend},{\{1,\dots,n\}}}$. We then align these two sets of point clouds using Chamfer loss to optimize the global transformation (rotation, scale, and translation) over 2000 iterations. To prevent the alignment from becoming trapped in local minima, such as an object being flipped along its vertical axis, we manually flip the point clouds along the axis starting at the 1000th iteration. We then compare the Chamfer loss before and after flipping to select the best alignment. After aligning the point cloud sets, we optimize the camera pose for 500 iterations using MSE loss between the 3D point clouds of $\mathrm{PC}_{\text{ref}}$ and the 2D pixel positions in $V_{\text{ref},t}$.
@@
We enable optimization of the negative condition embedding in TRELLIS. At each inference step of the generative model (conditioned on this embedding), we decode the output into a mesh. By computing the visual alignment between the rendering view of the reconstructed mesh and Gaussian Splats, we backpropagate the gradient through differentiable rendering to the negative condition embedding in TRELLIS. We use a combination of DreamSim, LPIPS, and MSE losses (each with a weight of 1), along with an MSE regularization term weighted at 0.2 to compare with the embedding’s initial value, preventing distortion of the reconstructed result. This optimization is applied only during the latter part of the flow model ($0.6\geq t\geq 0$, where $t$ represents the timestep of the flow model). We begin the optimization at iteration 5 and gradually increase by 1 every 5 timesteps of the flow model.
@@
Total Mesh Gen Camera Pose Search
DUSt3R PSO SGD
57.7s 1.5s 13.9s 11.1s 0.5s
Mesh Refine Mesh Regist Texture Map Optim Mesh Interp 4D Asset Convert
14.1s 15.6s 0.3s 0.6s 0.1s
Table 4: Detailed average runtime per frame for each component of our framework.
@@
All experiments in this paper were conducted on a single A100 GPU. Table 4 provides detailed runtime information for each framework component as a supplement to Table 3 in the main paper.
@@ Appendix B Ablation Studies
B.1 Quantitative Ablation Studies
Ablation Study on Different Base 3D Generation Models.
In Table 5, we replace our base 3D generator, TRELLIS, with several contemporary models, including Hunyuan3D-2.0 [67], TripoSG [21], and CraftsMan3D [20]. The results demonstrate that using more advanced 3D generators (such as Hunyuan3D-2.0 and TripoSG) leads to corresponding improvements in performance, illustrating the extensibility of our method alongside ongoing advancements in 3D generation techniques.
[…]
Simple Naïve TRELLIS 0.8905 0.1597 1342.66 0.1282
Naïve CraftsMan3D 0.9177 0.1702 1185.47 0.0924
Naïve TripoSG 0.9259 0.1534 973.16 0.0657
Naïve Hunyuan3D 2.0 0.9185 0.1471 851.01 0.0647
V2M4 (TRELLIS) 0.9259 0.1017 825.59 0.0688
V2M4 (CraftsMan3D) 0.9212 0.1051 960.63 0.0772
V2M4 (TripoSG) 0.9286 0.0971 758.25 0.0652
V2M4 (Hunyuan3D 2.0) 0.9286 0.0885 691.48 0.0634
Complex Naïve TRELLIS 0.8887 0.1265 1216.19 0.1492
Naïve CraftsMan3D 0.9060 0.1350 1162.00 0.1027
Naïve TripoSG 0.9201 0.1404 1023.88 0.0891
Naïve Hunyuan3D 2.0 0.9097 0.1185 1022.81 0.0962
V2M4 (TRELLIS) 0.9008 0.0747 666.04 0.1220
V2M4 (CraftsMan3D) 0.9268 0.0709 558.34 0.1046
V2M4 (TripoSG) 0.9359 0.0608 433.38 0.0873
V2M4 (Hunyuan3D 2.0) 0.9192 0.0605 415.53 0.0990
All Naïve TRELLIS 0.8891 0.1352 1014.45 0.1438
Naïve CraftsMan3D 0.9098 0.1446 970.24 0.0999
Naïve TripoSG 0.9222 0.1436 859.04 0.0833
Naïve Hunyuan3D 2.0 0.9144 0.1284 796.98 0.0853
V2M4 (TRELLIS) 0.9073 0.0817 576.73 0.1082
V2M4 (CraftsMan3D) 0.9270 0.0795 524.00 0.0977
V2M4 (TripoSG) 0.9359 0.0698 401.71 0.0818
V2M4 (Hunyuan3D 2.0) 0.9224 0.0674 377.14 0.0901
Table 5: Impact of different base 3D generators on the overall performance.
[…]
TRELLIS Final 0.9259 0.1017 825.59 0.0688
w/o PSO 0.8855 0.1158 1090.86 0.1184
w/o DUSt3R 0.9236 0.1046 756.29 0.0713
w/o SGD 0.9146 0.1269 1086.14 0.0861
w/o Mesh Refinement 0.9028 0.1103 1023.60 0.1032
Replace DUSt3R with VGGT 0.9230 0.1006 821.22 0.0631
CraftsMan3D Replace DUSt3R with VGGT 0.9207 0.1046 887.53 0.0753
TripoSG Replace DUSt3R with VGGT 0.9250 0.0960 730.13 0.0757
Hunyuan3D-2.0 Replace DUSt3R with VGGT 0.9279 0.0879 727.24 0.0636
Table 6: Ablation study about the impact of key components in our framework.
[…]
DreamSim
(Camera Search)
0.1 0.9176 0.1058 824.20 0.0784
1 0.9259 0.1017 825.59 0.0688
10 0.9234 0.1037 853.18 0.0726
DreamSim
(Mesh Refinement)
0.1 0.9129 0.1054 876.41 0.0911
1 0.9259 0.1017 825.59 0.0688
10 0.9226 0.1078 932.72 0.0750
ARAP 0.1 0.9200 0.1046 814.31 0.0727
1 0.9259 0.1017 825.59 0.0688
10 0.9258 0.1058 928.54 0.0744
Table 7: Impact of varying DreamSim and ARAP loss weights.
[…]
Original 0.9259 0.1017 825.59 0.0688
0.1$\times$ 0.9289 0.1022 942.99 0.0686
1$\times$ 0.9259 0.1077 947.57 0.0771
10$\times$ 0.8016 0.1694 2410.03 0.2577
Table 8: Noise impact.
@@ B.1 Quantitative Ablation Studies
Ablation of Components and Framework Complexity Analysis.
Table 6 presents an ablation study of the key components in our framework, confirming the effectiveness of each procedure. We also evaluate the impact of replacing the default dense stereo reconstruction model DUSt3R with the more advanced VGGT [48], demonstrating additional performance improvements.
Regarding theoretical complexity, our framework exhibits linear time complexity $\mathcal{O}(T)$ with respect to the number of frames $T$, while maintaining almost constant memory usage $\mathcal{O}(1)$.
Quantitative Parameter Tuning for DreamSim and ARAP.
As outlined in Appendix A, during the optimization process, we assign only a loss weight factor for DreamSim and ARAP losses, which are set to 1 by default. Table 7 illustrates the performance impact of varying these parameters.
Robustness to Base Mesh Quality.
As discussed in the limitations section of the main paper, the quality of the base mesh significantly impacts the results. Table 5 shows that higher-quality meshes produced by advanced 3D generators (e.g., Hunyuan3D-2.0 and TripoSG) yield improved robustness. Additionally, Table 8 provides a detailed analysis of robustness by evaluating the effects of injecting varying levels of Gaussian noise during mesh generation.
B.2 Qualitative Ablation Studies